OpenStudy (anonymous):

the inlet pipe on a water tank can fill the tank in 8 hrs. when the tank was full, both the inlet pipe and the drainpipe were accidentally opened. twenty four hours later, the tank was empty. how many hours would it take to empty a full tank if only the drain were open

OpenStudy (anonymous):

I believe we must assume that the take empties/fills at constant rates.

OpenStudy (anonymous):

Let \(f\) be the fill time and \(d\) be t he drain time.

OpenStudy (anonymous):

well, I mean fill rate and drain rate

OpenStudy (anonymous):

We know that the size of the tank is \(8f\) and it is \(24(f-d)\). So you are trying to find \(t\) such that \(td\) is the side of the tank.

OpenStudy (anonymous):

You want to find \(t\). You have two equations: \[ 24(f-d)=8f\\ td = 8f \]

OpenStudy (anonymous):

@KawasumiKimiko Do you know how to solve this?

OpenStudy (anonymous):

quadratic?

OpenStudy (anonymous):

No, you can use substitution

OpenStudy (anonymous):

Okay, can you isolate \(d\) in the first equation?

OpenStudy (anonymous):

distribute 24?

OpenStudy (anonymous):

Yeah, try that first

OpenStudy (anonymous):

24f-24d=8f

OpenStudy (anonymous):

Okay, you are close, but can you isolate \(d\)?

OpenStudy (anonymous):

uhm. subtract 24f from both sides & then divide both sides by -24?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so d=16f/-24

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Now plug that in for \(d\) in the second equation

OpenStudy (anonymous):

You'll be able to solve for \(t\), because the \(d\) cancels.

OpenStudy (anonymous):

Can you do it?

OpenStudy (ganpat):

I might have another approach to the problem, if the answer is right :p Tank filled in an hour = 1/8 Tank emptied in an hour = 1/B... assuming B is total drain time Total time for the tank to completely fill or drain when both the pipes are left open = 24 hrs 1/8 + 1/B = 1/24 1/B = 32/192 B = 6 hrs..